OpenStudy (anonymous):

The ratio of factorial of a number x to the square of the factorial of other number,which when increased by 50% gives required number,is 1.25. Find the number x?

OpenStudy (anonymous):

let the another number be y then, y+50% of y = x 1.5y=x => y= x/1.5 also, x!/(y!)^2 =1.25 x! / [ (x/1.5)!]^2 =1.25

OpenStudy (anonymous):

since factorial is only of integer numbers so, x = 1.5*2*n , where n is any natural number

OpenStudy (anonymous):

so, x! / [ (x/1.5)!]^2 =1.25 => (3n)! / [ (2n)!]^2=1.25 where x=3n

OpenStudy (anonymous):

then?

OpenStudy (anonymous):

now, when n=1, (3!)/[(2!)]^2 = 1.5 which is not equal to 1.25 when n=2 , (6!)/[(4!)]^2 = 1.25 hence , x=3n=6